### v1

# Beyond Needle(s) in the Embodied Haystack: Environment, Architecture, and Training Considerations for Long Context Reasoning

Abstract
We introduce $\infty$-Thor, a new framework for long-horizon embodied tasks that advances long-context understanding in embodied AI. $\infty$-Thor provides: (1) a generation framework for synthesizing scalable, reproducible, and unlimited long-horizon trajectories; (2) a novel embodied QA task, Needle(s) in the Embodied Haystack, where multiple scattered clues across extended trajectories test agents’ long-context reasoning ability; and (3) a long-horizon dataset and benchmark suite featuring complex tasks that span hundreds of environment steps, each paired with ground-truth action sequences. To enable this capability, we explore architectural adaptations, including interleaved Goal-State-Action modeling, context extension techniques, and Context Parallelism, to equip LLM-based agents for extreme long-context reasoning and interaction. Experimental results and analyses highlight the challenges posed by our benchmark and provide insights into training strategies and model behaviors under long-horizon conditions. Our work provides a foundation for the next generation of embodied AI systems capable of robust, long-term reasoning and planning. The datasets and code can be found at pearls-lab.github.io/infini-thor.
1 Introduction
Real-world embodied reasoning is a sequential decision-making problem requiring long-horizon planning, where task success depends on both memorizing and reasoning over multiple events that occur far apart in time. Using large pre-trained vision-language-action (VLA) models as policies for such tasks requires surpassing the key challenge of long-context reasoning. We seek to answer questions pertaining to what design choices matter in terms of environments, model architectures, and training methods when using VLA models for long-horizon embodied tasks. To this end, we develop a new framework for long-horizon tasks designed to push the boundaries of long-context understanding in embodied AI.
Figure 1: Example of the trajectory and a long-horizon embodied task generated from $\infty$-Thor. The final goal (“Put the tomato on the counter top” at t=670) requires recalling both the tomato (seen at t=17) and the counter (seen at t=560) to solved the long-horizon task. Context size refers to the input token length when converting the trajectory into the LLM input space.
We introduce $\infty$-Thor, a new framework for generation, training, and evaluation of long-horizon embodied tasks. Our benchmark uniquely features tasks with a synthetic final goal, which involves multiple objects that appear at distant time steps, requiring multi-step reasoning across over hundreds of steps. Figure 1 illustrates an example: the agent observes the tomato at an early step (t=17) and the counter top much later (t=560). Then, the final task is given at t=670, which requires the agent to place the tomato on the counter top. This setup highlights the challenge of long-horizon dependency, where key objects and locations must be remembered and acted upon after hundreds of steps.
This long-horizon setup introduces a new challenging task, Needle(s) in the Embodied Haystack (NiEH). Unlike the standard Needle in a Haystack task liu-etal-2024-lost , which focuses on recalling a single clue in text, NiEH poses two main challenges: (1) multiple scattered clues (Needles) and (2) multi-modal inputs that combine visual and linguistic observations from the environment (Embodiment). This task is designed to evaluate the agent’s ability to recall and reason about previously encountered environmental details, such as identifying objects and recalling performed actions.
Going beyond static evaluations such as NiEH, $\infty$-Thor also provides an interactive evaluation, allowing agents to execute policies and complete long-horizon tasks within a dynamic environment. To support this, we release a trajectory dataset for training, with episodes over 400 steps in the training set and more than 600 steps in the dev and test sets. These trajectories can be used for imitation learning, and our experiments show that access to longer context during training leads to significant performance gains, highlighting the importance of our dataset for long-context embodied reasoning.
We further investigate various architectural considerations for embodied agents to operate under extreme sequence lengths. We show that interleaved Goal-State-Action modeling—a multimodal, goal-conditioned VLA architecture that jointly models interleaved sequences of goals, states, and actions using a LLM backbone is the most practical approach for this class of problems. Moreover, since standard LLMs are constrained by fixed context windows and cannot natively handle inputs exceeding 1M tokens, we explore long-context extension techniques such as rotary embedding scaling and positional interpolation chen2023extendingcontextwindowlarge ; ding2024longropeextendingllmcontext ; peng2024yarn . Lastly, we demonstrate how to further strengthen long-context reasoning by fine-tuning the model on extended-context inputs using Context Parallelism, a parallel training strategy that allows efficient scaling to very long sequences.
We provide comprehensive experiments and analyses, demonstrating both the challenges posed by our benchmark and the behavior of baseline models under long-horizon settings. We investigate a range of training considerations, including different configurations for fine-tuning and long-context adaptation, and evaluate their impact on model performance.
Our contributions are summarized as follows:
  • We introduce $\infty$-Thor, a new framework for generating, training, and evaluating long-horizon embodied tasks, featuring synthetic final goals that require multi-step reasoning across hundreds of steps.
  • We propose a novel embodied QA task, Needle(s) in the Embodied Haystack, requiring agents to recall and reason over multiple scattered clues across extended trajectories.
  • We release a large-scale trajectory dataset and an interactive evaluation environment to support both offline imitation learning and online policy execution in long-horizon settings.
  • We describe architectural adaptations including interleaved Goal-State-Action modeling, long-context extension and Context Parallelism, tailored for interactive embodied reasoning.
  • We present empirical results and analyses, providing insights to the current capabilities and limitations of embodied AI systems on long-horizon tasks.
2 Related Work
Table 1: Comparison of benchmarks. We use Short (< 50 steps), Medium (50–300 steps), and Long (> 300 steps) to describe task horizon, reflecting the approximate number of environment steps required to complete a task in each benchmark. Single/Multi in the QA set column denotes single- and multi-evidence question type.
Benchmark Task Interaction Dataset QA set
/Platform Horizon w/ env modality # steps GT actions single multi
ProcTHOR deitke2022procthorlargescaleembodiedai
MineDojo fan2022minedojobuildingopenendedembodied Long
Habitat 3.0 puig2023habitat30cohabitathumans Long
VirtualHome puig2018virtualhomesimulatinghouseholdactivities Short multi 11.6
ALFRED shridhar2020alfred Medium multi 50
ALFWorld ALFWorld20 Medium text 50
BEHAVIOR-100 srivastava2021behaviorbenchmarkeverydayhousehold Med/Long
BALROG author2024BALROG Long
EQA embodiedqa
MM-EGO ye2025mmego
$\infty$-Thor Long multi 600+
Long-horizon Planning in Virtual Environments. AI2THOR kolve2017ai2 provides interactive indoor environments widely used for embodied reasoning research, while ProcTHOR deitke2022procthorlargescaleembodiedai extends these capabilities by procedurally generating scalable environments, potentially facilitating longer trajectories. MineDojo fan2022minedojobuildingopenendedembodied offers an open-ended platform within Minecraft, explicitly geared toward tasks requiring extensive long-term planning. Additionally, platforms such as VirtualHome puig2018virtualhomesimulatinghouseholdactivities and Habitat 3.0 puig2023habitat30cohabitathumans have demonstrated suitability for tasks involving long-term interactions and complex activity sequences. However, all of these platforms only provide environments and do not include standardized datasets or benchmark suites to support training and evaluation for long-horizon embodied tasks.
Embodied QA and Multimodal Needle in the Haystack Tasks. Embodied QA tasks, such as EmbodiedQA embodiedqa and MM-EGO ye2025mmego , require agents to answer questions grounded in visual observations, often demanding spatial and temporal reasoning over the agent’s trajectory. While these benchmarks emphasize multimodal understanding, they do not involve active interaction with the environment during evaluation. Another related area to our NiEH task is the multimodal Needle in a Haystack (NiH) problem. Traditional NiH tasks primarily assessed textual recall within long-context inputs liu-etal-2024-lost , while recent multimodal variants extend this idea by incorporating visual components wang2024needlemultimodalhaystack ; wang-etal-2025-multimodal . However, these works operate on relatively short context windows (typically up to 72K tokens) and do not require embodied reasoning or temporal dependencies across a dynamic trajectory.
Datasets and Benchmarks for Long-horizon Embodied Tasks. Recent efforts have pushed toward long-horizon embodied tasks, where agents must complete multi-step goals with extended temporal dependencies. While ALFRED shridhar2020alfred and ALFWorld ALFWorld20 introduced multi-step instruction-following tasks with action annotations and textual grounding, their task horizons are relatively short, typically under 50 steps. BEHAVIOR-100 srivastava2021behaviorbenchmarkeverydayhousehold evaluates agent generalization on household activities, some of which require prolonged engagement, but mainly focus on single task. BALROG author2024BALROG is a benchmark for testing the agentic capabilities of long-context LLMs, but its scope is limited to games.
Long-context Benchmarks. Outside embodied AI, general benchmarks have addressed challenges in long-context reasoning. Benchmarks, such as LongBench bai-etal-2024-longbench and RULER hsieh2024rulerwhatsrealcontext , focus on retrieval or summarization tasks. GSM-$\infty$ zhou2025gsminfinitellmsbehaveinfinitely extends GSM-8K cobbe2021trainingverifierssolvemath to assess mathematical reasoning over extremely long textual inputs.
3 $\infty$-Thor: An Environment for Generating, Training, and Evaluating Long-horizon Embodied Tasks
$\infty$-Thor features with a generation framework for synthesizing long trajectories to train and evaluate AI agents in long-horizon embodied tasks. We build $\infty$-Thor upon the AI2-THOR simulator kolve2017ai2 , an interactive 3D environment for embodied AI research that supports diverse scenes, objects, and agent actions. $\infty$-Thor enables the creation of trajectories with no length limit, and provides an evaluation setup where agents can interact dynamically with the environment during both training and testing. This supports both offline learning by producing large-scale datasets, and online learning through direct agent-environment interaction.
Each trajectory generated by $\infty$-Thor consists of multiple task goals, such as “Put a clean sponge on a metal rack” and “Pick up the apple and place it on the microwave”, requiring grounded understanding and action to achieve the goal. This enables the agent to explore and interact with the environment over extended episodes. At the end of each trajectory, the agent is assigned a synthetic long-horizon task that requires reasoning over entities encountered at distant time steps. For the example in Figure 1, the long-horizon task (Sub-goal #23) at step t=689, “Put the tomato on the counter top”, depends on observations made far earlier: the tomato at t=17 and the counter top at t=560. Our generation framework can generate unlimited tasks, the trajectories can be exceptionally long, exceeding 1M context tokens or beyond when the trajectory is processed with LLMs.
Successfully completing this task requires the agent to (1) memorize and integrate key environmental information over hundreds of steps, and (2) plan actions based on dependencies that are separated in time, demonstrating the need for long-context reasoning and robust spatio-temporal memory.
3.1 Static Evaluation: Needle(s) in the Embodied Haystack
(a) Needle in the Embodied Haystack: Single-evidence question types.
(b) Needles in the Embodied Haystack: Multi-evidence question types.
Figure 2: Example of N(s)iEH task and Ground-truth steps.
We first introduce a novel task in the form of a static evaluation: Needle(s) in the Embodied Haystack (NiEH). NiEH is designed to evaluate an agent’s ability to recall and reason about environmental states encountered throughout a trajectory. Unlike traditional embodied QA tasks that focus primarily on visual understanding of a single image, NiEH emphasizes reasoning about environmental changes over time, requiring agents to interpret and integrate sequences of multimodal observations.
Figure 2 presents examples of the two NiEH task types. In the single-evidence setting, a question is answerable based on a single observation step; in the multi-evidence setting, multiple temporally distant steps must be combined to answer the question. The NiEH testset includes diverse question types, such as binary (“yes” or “no”), “what”-, “where”-, and “how many”-style questions. These questions span a broad range of difficulty, from simple memory recall (similar to the Needle in a Haystack paradigm) to complex queries that requiring multi-step reasoning across temporally and spatially distributed evidence.
Testset Construction. We first replay the generated trajectories and collect the agent’s egocentric views, along with all objects that interact with the agent throughout the trajectories, such as objects that are picked up, moved, or even simply observed. Based on these interactions, we apply a set of rule-based templates to generate QA pairs, such as “Q. What object did you slice? A. {object name}” and “Q. Is {obeject name} on the desk? A. Yes/No”. Then, we sample questions based on the frequency to ensure diversity across object types, and annotate the GT answer steps using the replay logs.
After generating QA pairs and annotating the GT steps, we cross-validate the answerability of each question with GT images using four different multimodal LLMs: LLaVA-OneVision 7B li2024llava , Qwen2.5-VL 7B bai2025qwen25vltechnicalreport , Deepseek-VL 7B lu2024deepseekvlrealworldvisionlanguageunderstanding , and Pixtral 12B agrawal2024pixtral12b . Since these models are highly capable at standard visual QA, we filter out the questions that none of the four models successfully answer with GT images. At test time, the entire trajectory is treated as a Haystack and then cropped based on the GT image’s depth. Full details on templates, generation rules, and the validation scores of the four models are included in the Appendix.
Challenges in Needle(s) in the Embodied Haystack. The NiEH task introduces two key challenges for current models. First, many questions require reasoning over multiple temporally distant events. As shown in Figure 2(b), the agent moves a dish sponge from the garbage bin at $t=24$, then to the counter top, and later places it into a drawer at $t=751$. A question such as “Where was the dish sponge before you put it on the counter top?” requires the model to recall and chain together multiple actions and locations across hundreds of steps. Second, some questions demand aggregating sparse and temporally scattered evidence from long trajectories. In the second example in Figure 2(b), answering “How many times did you move the credit card?” requires the model to track and count all relevant actions occurring from the beginning to the end of the episode. These challenges highlight the need for models that can perform robust long-horizon reasoning across both time and modalities in complex embodied environments.
Table 2: Dataset Statistics
NiEH testset Single-clue Multi-clue
# of question-answer pair 829 474
Trajectory Train Dev Test
# trajectory 2,456 125 225
# avg/max sub-goals 12/18 16/24 18/33
# avg steps 405 613 627
# max steps 654 890 952
# avg token length 602K 880K 912K
# max token length 954K 1.2M 1.3M
3.2 Constructing Long-horizon Trajectories for Interactive Evaluations
With $\infty$-Thor’s generation framework, we can synthesize long-horizon trajectories to construct training, validation, and test sets for offline learning and evaluation. Our approach builds upon a planner-based method kolve2017ai2 , in which we sequentially concatenate multiple single-task demonstrations into a extended trajectory, while maintaining consistency in object states and agent interactions throughout.
To generate each trajectory, we first sample a task type from one of seven predefined templates (e.g., pick two objects and place, pick and place with movable receptacle). We then sample objects that are required to perform the task, such as items to be picked up or receptacles to be interacted with. Based on the sampled task and objects, we use a classical task planner that operates on PDDL-defined domains to generate ground-truth action sequences. These plans are executed in a simulator, and only successful rollouts are retained. We then concatenate these successful demonstrations to construct long-horizon sequences that span hundreds of steps. For the final goal, the involved objects are sampled exclusively from those seen during the early 20% and the final 20% of the trajectory. This enforces a long-term temporal dependency between two objects that must be jointly referenced to complete the final task. Through this procedure, we generate 2,456/125/225 trajectories for the training, validation, and test sets, respectively. Details on task types and a pseudo-algorithm for the generation process are provided in the Appendix.
4 Architectures for Long-Horizon Vision-Language-Action Models
Figure 3: Agent–environment interaction through interleaved Goal-State-Action modeling.
Embodied agents must effectively interact with complex, dynamic environments, necessitating capabilities to interpret multimodal inputs (vision, language) and produce coherent sequences of actions. Developing such Vision-Language-Action (VLA) models is particularly challenging due to the need for seamless integration of perceptual understanding, linguistic reasoning, and action prediction. Existing VLA models either use separate encoders for vision, language, and action modules shridhar2020alfred , or focus on short-horizon, low-level tasks such as robot arm manipulation in constrained environments brohan2023rt1roboticstransformerrealworld ; kim24openvla , where decisions depend only on the most recent observation and a single instruction. While recent multimodal LLMs like LLaVA NEURIPS2023_6dcf277e , MiniGPT-4 zhu2023minigpt , and Llama 3.2 meta2024llama32 show strong multimodal reasoning abilities, they primarily operate on static inputs (e.g., single or few images) and lack the dynamic interactivity and memory needed for long-horizon embodied tasks involving continuous vision-language-action sequences. Moreover, many state-of-the-art models are only accessible via proprietary APIs, making them impractical for real-time, controllable embodied settings and managing long-term memory states.
We explore the potential of a multimodal LLM as a unified model for VLA modeling, utilizing an interleaved input structure of goal, state (visual observations), and action tokens, as illustrated in Figure 3. This interleaved multimodal input allows the model to process vision, language, and actions concurrently, thereby facilitating more coherent, real-time interaction modeling. Specifically, our goal-conditioned agent uses a multimodal LLM backbone trained to predict subsequent actions autoregressively, conditioned on sequences of goal and state tokens. At each timestep $t$, the environment provides a new visual observation $s_{t}$, which is encoded as state tokens and appended to the existing token stream. The model then autoregressively predicts the next action $a_{t}$ conditioned on the full history of goals, states, and previously taken actions. This action is executed in the environment (e.g., “Pick up the book”), which leads to the next observation $s_{t+1}$, continuing the perception-action loop. This interactive sequence is repeated over hundreds of steps, allowing the model to reason over temporally distant information while maintaining grounded behavior in dynamic settings. By leveraging this interleaved Goal–State–Action modeling, our architecture supports coherent decision-making across long-horizon embodied tasks.
Context Extension. Given the limitations in context length of most LLMs, using off-the-shelf models is insufficient for processing long inputs such as those exceeding 1M tokens. We explore various long-context extension techniques that allows the model to generalize to longer input sequences without retraining from scratch. Specifically, we consider: Linear Interpolationchen2023extendingcontextwindowlarge : Rescales input positions to fit within the pretrained RoPE range by linearly interpolating positional indices; Dynamic Scalingchen2023extendingcontextwindowlarge : Adapts RoPE frequencies at runtime based on the input sequence length, using a linear rescaling to maintain consistent positional encoding behavior across varying lengths; YaRNpeng2024yarn : Dynamically interpolates attention frequencies during inference, balancing between pretrained and extrapolated positional regimes; LongRoPEding2024longropeextendingllmcontext : Augments RoPE with specially designed extrapolation functions, enabling robust generalization to long sequences without degrading attention quality. We apply these techniques during fine-tuning, at inference time, or both.
Context Parallelism. To further enhance the model’s ability to reason over long contexts, it is crucial to fine-tune on extended context inputs. However, the quadratic complexity of the attention mechanism makes it computationally infeasible to train directly over long sequences. To address this challenge, we employ Context Parallelism, a parallel training technique designed for efficient long-context modeling.
Context Parallelism leverages Ring Attention liu2023ring , a novel parallel implementation of the attention layer. In Ring Attention, key-value (KV) shards are cyclically shuffled across devices, and partial attention scores are computed iteratively. This process is repeated until all KV shards have been incorporated on each device, ensuring complete attention coverage without the full memory cost of standard attention. By combining Context Parallelism with our dataset of extended long-context inputs, we are able to scale fine-tuning to substantially longer sequences, unlocking improved long-horizon reasoning capabilities.
5 Experiments
5.1 Static Evaluation: Needle(s) in the Embodied Haystack
We first evaluate model performance on the Needle in the Embodied Haystack (NiEH) and Needles in the Embodied Haystack (NsiEH) tasks, which test an agent’s ability to retrieve and reason over sparse evidence scattered throughout long embodied trajectories.
Building a Embodied Haystack. Unlike the traditional Needle in the Haystack setup, which inserts a target sentence into a long text corpus like a book, we use the entire embodied trajectory as the input context. To simulate different reasoning depths, we crop the input sequence either from the beginning or the end based on the GT image’s position. In the NsiEH task, where multiple evidences are scattered throughout the trajectory, we fill the context with intermediate steps in between the GT steps keeping their temporal order.
Results. Figure 4 presents the performance of LLaVA-OneVision 7B li2024llava model across various context extension methods. Linear Interpolation, Dynamic Scaling, and LongRoPE scaling all struggled with very long contexts beyond 128K tokens (the results of Linear Interpolation are excluded from Figure since it fails at all examples). YaRN consistently outperformed other methods across both NiEH and NsiEH, successfully answering questions at context lengths exceeding 384K tokens, likely due to its architectural alignment with LLaVA-OneVision’s Qwen2 LM backbone, which employs RoPE and YaRN scaling during pretraining. YaRN performed best at moderate scaling factors (e.g., x4), however, further scaling to x8 and x16 did not yield additional gains. In particular, x16 slightly improved performance in the 256K–384K token range but led to degradation notably at shorter context sizes (<64K), suggesting that excessive scaling may introduce instability and negatively impact performance. Overall, all methods fail beyond 512K tokens. We expect these trends to persist as models are pushed to even larger context windows, highlighting the need for more advanced methods to effectively handle extremely long-context scenarios in the future.
Single vs Multi-evidence Reasoning. Comparing NiEH to the more challenging NsiEH task, we observe a significant performance drop in the multi-evidence setting. This is especially pronounced for mid-depth questions involving sparse or distant evidence (e.g., “Where was the Mug before you put it on the CounterTop?”) or questions requiring the aggregation of multiple clues (e.g., “How many times did you move the Apple?”), as shown in Figure 2(b). These results demonstrate that our NiEH and NsiEH tasks pose a substantial challenge to current long-context models and success requires both fine-grained temporal memory and multi-evidence reasoning across extended interactions.
(a) Results of Needle in the Embodied Haystack (NiEH).
(b) Results of Needles in the Embodied Haystack (NsiEH).
Figure 4: Results of Needle(s) in the Embodied Haystack with the LLaVA-OneVision 7B model. The white dashed line denotes the maximum input context length of the model. Context Parallelism is applied to all experiments with the context size over 384K.
5.2 Interactive Evaluation in $\infty$-Thor
We conduct an interactive online evaluation using the AI2THOR simulator to measure agent performance on our long-horizon test set. Our experiments analyze reward accumulation across various context extension methods, different scaling factors, and multiple fine-tuning context lengths: 32K, 64K, and 130K.
Training. We fine-tune the LLaVA-OneVision 7B model on our training set while freezing the vision encoder. We used 8 H100 GPUs with both tensor parallelism shoeybi2020megatronlmtrainingmultibillionparameter and pipeline parallelism huang2019gpipeefficienttraininggiant for the 32K context size, while Context Parallelism is employed for training on larger context sizes (64K and 130K). Additional training specifics are available in the Appendix.
Figure 5: Agent’s reward across different experimental configurations. We compare (a) context extension methods at fixed scaling (x4), (b) varying YaRN scaling factors, (c) fine-tuning with different context lengths using Context Parallelism, and (d-e) combinations of scaling during both training and inference. (f) summarizes the most effective strategies, highlighting that exposure to longer contexts during training significantly improves performance. Non-planner models cannot generate valid actions after around 250 steps ($\approx$376K in context size).
Plan-level Evaluation. We evaluate agent performance based on reward accumulation given previous states and actions. Evaluation is performed at the plan-level, where a plan represents a short sequence of actions aimed at a specific intermediate goal. For example, a “Go to location” plan comprises actions like Move Ahead, Rotate Right/Left, and Look Up/Down. Each trajectory is thus composed of multiple sequential plans. At each step, agents are presented with a task goal alongside the previous state-action sequences, predicting subsequent actions and interacting continuously with the environment until the plan completion. Further evaluation details are provided in the Appendix.
Results and Discussion. Figure 5 illustrates the comparative results across six experimental configurations. The Planner trajectory serves as the performance upper-bound. We address the following key questions through our experimental results:
Q. Which context extension methods perform best? Figure 5(a) compares different context extension methods at a fixed scaling factor of x4. Similar to the NiEH results, YaRN consistently achieves the highest performance showing very close performance to Planner.
Q. Does further scaling enhance performance? Figure 5(b) explores YaRN scaling at different scaling factors (x4, x8, and x16). Interestingly, increasing the scaling factor beyond x4 does not significantly improve performance, indicating a diminishing return for larger scaling factors.
Q. Is fine-tuning on a dataset with long trajectories effective? Figure 5(c) demonstrates the effectiveness of fine-tuning with Context Parallelism, enabling scaling of context lengths up to 64K and 130K tokens. At a 130K context size, the model can learn sequences comprising approximately 86 steps, substantially longer compared to only 22 steps with a 32K context size. This shows that exposure to longer context during training significantly enhances model performance, suggesting that incorporating more long-horizon data by $\infty$-Thor could further improve model capabilities. We note that context extension methods were not applied in this experiment.
Q. Does combining context extension methods during both training and inference provide additional benefits? Figures 5(d) and 5(e) illustrate experiments with scaling applied at both training and evaluation stages. Results indicate that additional scaling at evaluation after fine-tuning with scaled RoPE provides no further performance improvement and may degrade performance at shorter context lengths ($\leq$300K tokens).
Based on these observations, we can conclude that fine-tuning strategies are most effective when long-trajectory datasets are available. In the absence of extensive training data, employing YaRN scaling at x4 yields performance comparable to the Planner upper-bound, particularly within context lengths under 200K tokens (Figure 5(f)).
6 Conclusion
We presented $\infty$-Thor, a new framework for long-horizon embodied tasks designed to advance long-context understanding in embodied AI. Our framework enables scalable synthesis of long, complex trajectories paired with GT action sequences, and supports both offline training and online interaction with the environment. As part of this framework, we introduced a novel embodied QA benchmark, Needle(s) in the Embodied Haystack, that challenges agents to reason over sparse, temporally distant visual evidence embedded within extended trajectories. To equip models for this setting, we explored architectural adaptations including interleaved Goal–State–Action modeling, context extension techniques such as YaRN and LongRoPE, and efficient fine-tuning via Context Parallelism. Our experiments demonstrate that exposure to longer contexts during training significantly improves model performance, and the limitation of existing context extension techniques struggle with long-context reasoning. We hope our framework and benchmark encourage further research into models capable of robust long-horizon reasoning under realistic, interactive environments.
Appendix A Dataset Construction
A.1 Building the Needle(s) in the Embodied Haystack Benchmark
We construct the Needle(s) in the Embodied Haystack benchmark in three stages: 1) Trajectory Replay and Metadata Collection; 2) Rule-Based QA Generation; and 3) Cross-validation with Multimodal LLMs. The following sections provide detailed descriptions of each step.
A.1.1 Trajectory Replay and Metadata Collection
We first replay 225 test trajectories generated by $\infty$-Thor, logging both visual observations (agent’s egocentric views) and structured metadata at each timestep. For every step, we store the list of visible objects, agent-inventory items, openable containers, and their contents from the simulator. This produces a fine-grained interaction log that captures grounded scene dynamics over time.
An example of the collected metadata at a single timestep is shown below:
Example of metadata entry
    "img_idx": 2,
    "img_filename": "000000002.png",
    "step": 1,
    "object_log": {
        "visible": ["Shelf", "Vase", "Book"],
        "pickupable": ["Vase", "Book"],
        "isOpen": [],
        "inven_obj": [],
        "receptacles": ["Shelf"],
        "recep_objs": {
            "Shelf": ["Vase", "Book"]
Each metadata entry corresponds to a low-level action step and provides the semantic state of the scene, enabling the construction of temporally grounded QA instances in later stages.
A.1.2 Rule-Based QA Generation
To construct the QA set, we apply rule-based generation templates to each trajectory using its sequence of low-level actions and associated metadata. The QA generation process involves parsing the agent’s interactions with objects, containers, and the environment, and applying a set of handcrafted rules to synthesize grounded questions.
Our QA generation logic covers a diverse range of question types, including object presence, object state, location tracking, slicing actions, container content reasoning, and action counting. For instance, if an object is seen for the first time at a particular step, a presence question such as “Is there any apple in this room?” is generated. Similarly, after a PutObject action, location-based questions like “Where was the apple before you put it to the microwave?” are produced. When slicing actions happen, we create questions about the object being sliced and other nearby items (e.g., “What objects were in the Fridge when you sliced the apple?”). Then, we sample questions based on the frequency to ensure diversity across object types, and annotate the GT answer steps using the replay logs. Table 3 summarizes the types of questions generated, and corresponding trigger conditions and example templates.
Table 3: QA types, trigger conditions, and corresponding question templates used in rule-based generation.
QA Type Trigger Condition Example Template(s)
object presence (Yes/No) object appears visibly in the trajectory Is there any {obj} in this room?
Have you seen a/an {obj}?
open state questions container marked as open in metadata Was {container} open?
object location tracing sequences of Pickup and PutObject actions Where was {obj} before you put it to {container}?
Where did you move the {obj} from the {container}?
Where is {obj} now?
slicing-based questions SliceObject action detected in trajectory What did you slice?
What objects were in/on the {container} when you slice the {obj}?
container content container visibility with non-empty contents What objects were in/on the {container}?
What object did you put in/on the {container}?
put action questions unique PutObject action for a container What object did you put in/on the {container}?
final object state final location of an object at episode end Is {obj} in/on the {container}?
What objects are in/on the {container}?
How many objects were in/on the {container}?
movement counting object picked up more than once How many times did you move {obj}?
A.1.3 Cross-validation with Multimodal LLMs
To ensure the answerability and clarity of the generated QA pairs, we perform cross-validation using four powerful multimodal LLMs: LLaVA-OneVision 7B [19], Qwen2.5-VL 7B [3], Deepseek-VL 7B [23], and Pixtral 12B [1]. Each model is prompted with the GT images corresponding to the annotated QA steps and asked to answer the associated questions. Given their strong performance on standard visual QA tasks, we use these models to assess whether a question can be correctly answered or not. We keep only the QA pairs that are correctly answered by at least one of the four models, and discard those that fail across all models. This helps improve dataset quality and filtering out ambiguous or visually ungroundable questions. Table 4 shows the accuracy of each model on the finalized QA set when evaluated with GT images. Notably, even with access to GT images, all models struggle with questions requiring reasoning over three or more evidence steps. To maintain the benchmark’s difficulty and support evaluation of more capable models in future, we manually inspect the multi-clue questions and include those that are answerable.
Table 4: QA accuracy (%) of multimodal LLMs on ground-truth images.
Model Size # of clues (GT steps) Total
1 2 $\geq$3
LLaVA-OneVision 7B 86.61 68.55 23.74 71.15
Qwen2.5-VL 7B 85.94 89.83 64.40 82.20
Deepseek-VL 7B 81.56 39.14 22.57 62.88
Pixtral 12B 91.34 39.60 58.56 76.25
A.2 Constructing Long-Horizon Trajectories
To synthesize long-horizon trajectories, we construct each trajectory by sequentially chaining successful sub-tasks sampled from a predefined set of task templates. This process is illustrated in Algorithm 1. We begin by sampling a task template from a fixed task pool, which includes goal types such as pick and place simple, pick two obj and place, and pick and place with movable recep. Each sampled template requires relevant objects in the scene (e.g., pickupable items, target receptacles), which are then used to define the goal for that task.
We use a classical task planner, which operates over PDDL-defined domains [29], to generate a low-level action sequence for the sampled goal, and simulate this plan in an interactive environment. If the rollout fails (e.g., due to collisions, object occlusions, or unreachable conditions), we discard the sequence and re-sample from the task pool. Otherwise, the successful rollout is retained and appended to the ongoing trajectory.
This sampling-execution loop is repeated until a long trajectory with a desired number of sub-goals is formed. The resulting synthetic long-horizon trajectory consists of multiple sub-goals concatenated into a continuous sequence. To induce long-term temporal dependencies, the final sub-task is constrained to involve only objects that appear in the early 20% and late 20% of the overall trajectory, requiring the agent to integrate temporally distant evidence to answer associated questions.
Algorithm 1 Construct Long-horizon Trajectory
1:Input: Task Pool $\mathcal{T}$, max sub goals $N$
2:Output: Long-horizon trajectory $\tau$
3:Initialize empty trajectory $\tau\leftarrow[]$
4:while len($\tau$) $<$ $N$ do
5:     Sample task template $g\sim\mathcal{T}$ and objects
6:     Plan action sequence $\pi_{g}$ by planner
7:     if Simulate($\pi_{g}$) is successful then
8:         Append $\pi_{g}$ to trajectory: $\tau\leftarrow\tau\mathbin{\|}\pi_{g}$
9:     else
10:         Discard and re-sample
11:     end if
12:end while
13:// Final sub-task with long-term object dependency
14:Sample $g_{\text{final}}\sim\mathcal{T}$ and objects in early 20% and late 20%
15:Plan and simulate $\pi_{\text{final}}$ using restricted objects
16:if Simulate($\pi_{\text{final}}$) is successful then
17:     Append $\pi_{\text{final}}$ to trajectory: $\tau\leftarrow\tau\mathbin{\|}\pi_{\text{final}}$
18:else
19:     Repeat sampling until success
20:end if
21:return $\tau$
Appendix B Training and Evaluation Details
B.1 Interactive Evaluation in $\infty$-Thor
Training. We fine-tune the LLaVA-OneVision 7B model on our training set while freezing the vision encoder. The model is trained using a next-action prediction objective, where only the action tokens are optimized, conditioned on the goal and state tokens. Table 5 summarizes the training specifications for different context lengths. For 32K training, we apply tensor parallelism with a degree of 4 and pipeline parallelism with a degree of 2, utilizing 8 H100 GPUs in total. Since pipeline parallelism requires the batch size to match the pipeline degree, we set the batch size to 2. For longer context lengths, we use context parallelism: 8-way for 64K (on 8 GPUs) and 16-way for 130K (on 16 GPUs). All models are fine-tuned for approximately 3 epochs with a learning rate of 1e-5, using the AdamW optimizer and a linear learning rate schedule with a 0.03 warmup ratio.
Table 5: Training specifications for different context lengths.
Context Length Parallelism # GPUs Training Time
32K Tensor (4) + Pipeline (2) 8 160 hrs
64K Context (8) 8 120 hrs
130K Context (16) 16 134 hrs
Plan-Level Evaluation.
We evaluate agent performance using a plan-level framework, where each plan corresponds to a short sequence of actions aimed at achieving a specific intermediate sub-goal (e.g., navigating to an object, placing an item). A trajectory is composed of multiple such plans, executed sequentially. For the interactive evaluation, the agent is presented with the current plan’s goal along with the history of previous GT states and actions. Using this context, the agent predicts the next action and interacts step-by-step with the environment. The interaction continues until the current plan is either successfully completed or terminated due to failure (e.g., collisions or deadlocks). After each plan, the context is reset to include the GT actions and states from the completed portion of the trajectory, and the agent proceeds to the next plan. This ensures that each plan is evaluated independently, conditioned only on the correct prior history. The agent’s performance is measured via cumulative reward across all plans in the trajectory. Pseudocode for this evaluation procedure is provided in Algorithm 2.
Algorithm 2 Plan-Level Evaluation
1:Input: Trajectory $T=\{P_{1},P_{2},\ldots,P_{N}\}$, Agent policy $\pi$, Environment $\mathcal{E}$
2:Initialize: Reward $R\leftarrow 0$
3:Initialize state and history with initial observation
4:for each plan $P_{i}$ in $T$ do
5:     Initialize context with GT actions up to $P_{i-1}$
6:     while not done and not failure do
7:         $a_{t}\leftarrow\pi(\text{context})$
8:         $s_{t+1},r_{t},\textit{done},\textit{failure}\leftarrow\mathcal{E}.\text{step}(a_{t})$
9:         Append $(a_{t},s_{t+1})$ to context
10:         $R\leftarrow R+r_{t}$
11:     end while
12:     if failure then
13:         Break evaluation
14:     end if
15:end for
16:return Total accumulated reward $R$
Appendix C Limitations
While $\infty$-Thor enables the generation of arbitrarily long trajectories, the diversity of environment layouts in AI2-THOR is inherently limited. This can lead to repetitive agent behaviors within certain scenes. For instance, compared to kitchen and living room scenes, bedroom scenes tend to involve fewer action types, mostly constrained to simple pick-and-place tasks within small spatial areas. Due to the limited scene size and low task diversity, we excluded bathroom scenes from our dev and test sets. In future work, we plan to integrate $\infty$-Thor with ProcTHOR [9], which supports procedurally generated environments, enabling a broader range of dynamic and diverse scene configurations.
Additionally, although the GT action sequences generated by the PDDL-based planner are sufficient for task completion, they are not guaranteed to be optimal. This may lead to agents learning suboptimal behaviors when trained solely on these demonstrations. Incorporating learning from exploration or reinforcement-based optimization could improve policy quality.
Finally, inference with LLMs in long-context settings remains a computational bottleneck, particularly as context lengths approach 1M tokens. Since our long-horizon tasks require access to information spread throughout the entire trajectory, full-context inference becomes increasingly expensive, even with Context Parallelism. One promising direction is to equip agents with memory systems that selectively keep relevant information from prior steps, allowing the model to reason without reprocessing the full context at each step. Other architectural solutions, such as sparse attention mechanisms [13, 16] or state-space models with linear-time inference [12], also hold potential for scalable long-context reasoning.